Simple Interest
7. Suppose you purchase a six-year muni bond for $6000. The face value of the bond when it matures is $7620. Find the APR.

Answer by jorel555(1290)

7620-6000=1620
1620/6=270
270/6000=.045
So the APR is 4.5%. ☺☺☺☺